OpenStudy (anonymous):

Hello! I have a question! How does this: sqrt of (x^2-8x+16+9/16x^2) simplify to this: sqrt of (25/16x^2-8x+16) ??? Where the heck did the 25 come from? Thanks!

OpenStudy (anonymous):

\[\sqrt{x^2-8x+16+(9/16)x^2} > \sqrt{(25/16)x^2-8x+16}\]

OpenStudy (anonymous):

Just wanted to make it clearer for everyone to see. How the heck does it get from that to that???

OpenStudy (anonymous):

x^2 + (9/16)x^2 = (16/16)x^2 + (9/16)x^2 = (16/16 + 9/16)x^2 = (25/16)x^2

OpenStudy (anonymous):

Oh my god...thank you so much! That was so stupid of me! Thank you!